OpenStudy (anonymous):

Let f(x) = x/abs((x^2)-9) Find the domain of f. Write an equation for each vertical/horizontal asymptote of graph of f. Is f odd, even, or neither? Justify your answer. find all values of x for which f is discontinuous and classify each discontinuity as removable or nonremovable.

OpenStudy (anonymous):

first f is odd since f(-x)=-f(x) f is discontinous if x^2=9 i.e., x=3or x=-3 I think it is a non removable discontinuity domain of f is R-{-3,3}

OpenStudy (anonymous):

x=3 and x=-3 are asymptodes i am not sure but i think y=0 will be an asymptode too.

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}\frac{x}{\sqrt{(x^2-9)^2}}=\lim_{x \rightarrow \infty}\frac{x}{\sqrt{x^4-18x^2+81}}\] \[=\lim_{x \rightarrow \infty}\frac{\frac{1}{\sqrt{x^4}}}{\frac{1}{\sqrt{x^4}}} \cdot \frac{x}{\sqrt{x^4-18x^2+81}}\] \[\sqrt{x^4}|=|x^2|=x^2 \text{ for all } x \text{ since } x^2 \ge 0\] \[\lim_{x \rightarrow \infty} \frac{\frac{x}{x^2}}{\sqrt{\frac{x^4-18x^2+81}{x^4}}}\] \[\frac{0}{\sqrt{1-0+0}}=0\] y=0 is the horizontal asymptote